# Imbalanced Data Clustering using Equilibrium K-Means

### v1

@@ Abstract
Imbalanced data, characterized by an unequal distribution of data points across different clusters, poses a challenge for traditional hard and fuzzy clustering algorithms, such as hard K-means (HKM, or Lloyd’s algorithm) and fuzzy K-means (FKM, or Bezdek’s algorithm). This paper introduces equilibrium K-means (EKM), a novel and simple K-means-type algorithm that alternates between just two steps, yielding significantly improved clustering results for imbalanced data by reducing the tendency of centroids to crowd together in the center of large clusters. We also present a unifying perspective for HKM, FKM, and EKM, showing they are essentially gradient descent algorithms with an explicit relationship to Newton’s method. EKM has the same time and space complexity as FKM but offers a clearer physical meaning for its membership definition. We illustrate the performance of EKM on two synthetic and ten real datasets, comparing it to various clustering algorithms, including HKM, FKM, maximum-entropy fuzzy clustering, two FKM variations designed for imbalanced data, and the Gaussian mixture model. The results demonstrate that EKM performs competitively on balanced data while significantly outperforming other techniques on imbalanced data. For high-dimensional data clustering, we demonstrate that a more discriminative representation can be obtained by mapping high-dimensional data via deep neural networks into a low-dimensional, EKM-friendly space. Deep clustering with EKM improves clustering accuracy by 35% on an imbalanced dataset derived from MNIST compared to deep clustering based on HKM.
@@
K-means, fuzzy K-means, equilibrium K-means, imbalanced clustering, deep clustering
@@ I Introduction
Imbalanced data refers to data that is unevenly distributed across different clusters, which is prevalent in medical diagnosis, fraud detection, and anomaly detection. Learning from imbalanced data has been considered a challenge and a valuable direction [1]. While there is a considerable amount of research on supervised learning from imbalanced data, e.g., [2, 3, 4], unsupervised learning has not been as thoroughly explored, due to the unknown imbalance ratio between clusters making the task more challenging [5].
Clustering is an important unsupervised learning task involving grouping data into clusters based on similarity. K-means (KM) is the most popular clustering technique, valued for its simplicity, scalability, and effectiveness with real datasets. It is also used as an initialization method for more advanced clustering techniques, such as the Gaussian mixture model [6, 7]. KM starts with an initial set of centroids (cluster centers) and iteratively refines them to increase cluster compactness. The hard KM (HKM, or Lloyd’s algorithm) [8, 9] and fuzzy KM (FKM, or Bezdek’s algorithm) [10] are the two most representative KM algorithms. HKM assigns a data point to only one cluster, while FKM allows a data point to belong to multiple clusters with varying degrees of membership. In recent, they are also widely employed in deep learning, jointly trained with deep neural networks (DNNs) for clustering high-dimensional data, such as images [11, 12, 13, 14].
Many FKM variations have been developed to enhance its performance in different aspects. For example, the possibilistic FKM (PFKM) [15] was proposed to address the sensitivity of FKM to noise and outliers. FKM-$\sigma$ [16] was proposed to improve FKM’s performance on data points with uneven densities or non-spherical shapes in individual clusters. Fuzzy local information K-means (FLICM) [17] was designed to promote FKM’s performance in image segmentation. For more information on other FKM variants, please refer to the review paper [18].
Figure 1: Clustering results of a highly imbalanced dataset. (a) Ground truth. The colors represent the labels of the data points. (b) Clustering by hard K-means. (c) Clustering by fuzzy K-means. (d) Clustering by the proposed equilibrium K-means. They are all two-step alternating algorithms that iteratively update centroids and clusters.
Clustering imbalanced data also poses a challenge for both HKM and FKM. This is due to an undesirable property known as the “uniform effect”, which causes centroids to crowd together in large clusters, resulting in clusters of equal size even when input data has varied cluster sizes. An example illustrating the uniform effect of HKM and FKM is given in Fig. 1.
@@ I-A Existing Efforts to Overcome Uniform Effect
To improve the poor performance of FKM on imbalanced data, Noordam et al. [19] developed a modified FKM called cluster-size insensitive FKM (csiFKM). This method assigns weight to adjust the contribution of data points according to the size of the clusters they belong to. Lin et al. [20] proposed a size-insensitive integrity-based FKM (siibFKM) to reduce the sensitivity of csiFKM to the distance between adjacent clusters by increasing the cluster’s integrity. However, weighting based on cluster size inadvertently increases the influence of outliers, making these algorithms sensitive to noise [18].
Another approach to tackle the challenge of imbalanced data clustering is to use a multiprototype mechanism. This method first groups data into multiple subclusters, ensuring that the data points in each subcluster are relatively balanced. The final clusters are obtained by merging adjacent subclusters. Liang et al. [21] introduced a multiprototype clustering algorithm that employs FKM to generate subclusters. Later, Lu et al. [5] proposed a self-adaptive multiprototype clustering algorithm that can automatically adjust the number of subclusters. However, multiprototype clustering algorithms involve a complex process and high complexity, with a time complexity of $O(N^{2})$, where $N$ is the number of data points in the dataset, making them computationally expensive for large datasets. We should additionally mention that Zeng et al. [22] recently proposed a soft multiprototype clustering algorithm with time complexity linear to $N$. However, the clustering process of this algorithm remains complex and is aimed at clustering high-dimensional and complex-structured data rather than imbalanced data.
@@ I-B Our Contributions
In this paper, we propose equilibrium K-means (EKM) for imbalanced data clustering, which avoids the limitations of the aforementioned techniques. Our contributions can be summarized in three main aspects.
First, we offer a new framework that unifies HKM, FKM, and EKM with three key advantages. 1. This framework can directly re-derive HKM and FKM from unconstrained optimization problems with differentiable objectives without applying Lagrangian multipliers, simplifying the previous derivation from constrained optimization problems and non-differentiable objectives. 2. The new objective functions of HKM and FKM can be compared in the same plot, which used to be unfeasible due to the old objective function of FKM having more optimization variables than that of HKM. Such comparison can facilitate a more comprehensive and fundamental view than existing comparisons (such as [23, 24]) based on clustering outcomes. 3. This new framework allows for the uniform study of KM-type algorithms, such as the convergence conditions provided in this paper. We also reveal an explicit relationship between these KM algorithms and Newton’s method.
Second, we present EKM, derived from the new framework. Similar to FKM, EKM is a two-step alternating algorithm between updating the weighted centroid and calculating the weight, but the membership defined in EKM has a clearer physical meaning than that defined in FKM. Besides, EKM prevents centroids crowd together by introducing appropriate repulsive forces between centroids, making it more effective than HKM and FKM on imbalanced data. In addition to having the same time and space complexity as FKM, EKM also supports batch learning, making it scalable for large datasets. We demonstrate EKM’s performance through numerical experiments on two synthetic and ten real datasets from various domains. Because of space limitations, quantified metrics of performance are presented in the supplemental material, but the scatter graphs of partial clustering outcomes that we present in the paper are sufficient to demonstrate the superiority of EKM.
Finally, we investigate the cooperation between EKM and DNNs for high-dimensional data clustering. We find that mapping high-dimensional data via DNNs to a low-dimensional EKM-friendly space results in a more discriminative representation than mapping to an HKM-friendly space. When applied to an imbalanced dataset derived from MNIST, joint learning of DNNs and EKM improves clustering accuracy by 35% compared to joint learning of DNNs and HKM.
@@ I-C Organization
We introduce the HKM, FKM, and a variation of FKM, the maximum-entropy fuzzy clustering (MEFC) in Section II. In Section III and Section IV, we introduce the proposed clustering framework and the proposed EKM. We demonstrate the performance of EKM on classic clustering tasks in Section V and deep clustering in Section VI. Finally, we conclude in Section VII.
@@ II-A The Hard K-Means Algorithm (Lloyd’s Algorithm)
KM clustering aims to partition $N$ data points into $K$ clusters, minimizing the sum of the variances within each cluster. Mathematically, the objective can be expressed as:
@@
where $\mathbb{S}_{k}$ represents the set of data points in the $k$-th cluster, $\mathbf{x}\in\mathbb{S}_{k}$ denotes a data point belonging to $\mathbb{S}_{k}$, $\bm{\mu}_{k}$ is the centroid of points in $\mathbb{S}_{k}$, expressed by
@@
$|\mathbb{S}_{k}|$ signifies the size of $\mathbb{S}_{k}$, and $\|\cdot\|_{2}$ is the $l_{2}$ norm. The optimization problem (1) is NP-hard [25], and Lloyd‘s algorithm is a popular heuristic algorithm for solving this problem. Given $K$ points ($\mathbf{c}_{1}^{(1)},\cdots,\mathbf{c}_{K}^{(1)})$ as initial centroids, Lloyd’s algorithm alternates between two steps:
@@ II-B The Uniform Effect of K-Means
The uniform effect refers to the propensity to generate clusters of equal size, which is implicitly implied in the objective (1) of KM. For simplicity, kindly consider a case with two clusters (i.e., $K=2$). Minimizing the objective function (1) is equivalent to maximizing the following objective function [26]:
@@ II-C The Fuzzy K-Means Algorithm (Bezdek’s Algorithm)
$\min_{\mathbf{c}_{1},\cdots,\mathbf{c}_{K},\{u_{kn}\}}\quad$ $\sum_{n=1}^{N}\sum_{k=1}^{K}(u_{kn})^{m}\|\mathbf{x}_{n}-\mathbf{c}_{k}\|_{2}^{2}$ (6)
@@
    Calculate the membership values of data points of being in the cluster:
@@
    Recalculate the centroid of each cluster by taking the mean of all data points, weighted by their membership to the cluster:
@@
The time complexity of one iteration of the above two steps is $O(NK^{2})$. The higher time complexity of FKM than that of HKM is due to the extra calculation of membership. A practical time consumption comparison can be found in [27].
@@ II-D Maximum-Entropy Fuzzy Clustering
$\min_{\mathbf{c}_{1},\cdots,\mathbf{c}_{K},\{u_{kn}\}}\quad$ $\eta\sum_{n=1}^{N}\sum_{k=1}^{K}u_{kn}\ln{u_{kn}}$ (9)
@@
$\sum_{k}u_{kn}=1\,\forall n,$
where $\eta\in(0,1)$ is a hyperparameter, controlling the transition from maximization of the entropy to the minimization of centroid-data distances. MEFC is similar to FKM but with a different membership:
$u^{(\tau)}_{kn}=\frac{\exp{(-\lambda\|\mathbf{x}_{n}-\mathbf{c}_{k}\|_{2}^{2})}}{\sum_{i=1}^{K}\exp{(-\lambda\|\mathbf{x}_{n}-\mathbf{c}_{i}\|_{2}^{2})}},$ (10)
@@ III-A Objective of Smooth K-Means
Here, we propose a novel framework called smooth K-means (SKM) and demonstrate that the three KM algorithms introduced in the previous section are special cases of SKM. Based on this framework, we further develop a new clustering algorithm, aiming to improve the degraded performance of KM algorithms on imbalanced data. Denote the squared Euclidean distance [^1] Other distance metrics, such as the absolute difference and the angle between points, can also be used to define $d_{kn}$. Considering the derivation process remains the same, we use the Euclidean distance in the paper for simplicity. between the $k$-th centroid and the $n$-th data point as
@@
and define the within-cluster sum of squares (WCSS) as the sum of squared Euclidean distances between data points and their nearest centroid, i.e.,
@@
where $h(d_{1n},\cdots,d_{Kn})$ is a smooth approximation to $\min(d_{1n},\cdots,d_{Kn})$ and is referred to as the smooth minimum function. Below we explain the significance of replacing the minimum function with a smooth minimum function from a statistical point of view.
III-B Statistical Interpretation of Smooth K-Means
Statistically, the centroids obtained by minimizing WCSS are maximum likelihood estimators (MLEs) of parameters in a specific type of Gaussian mixture model (GMM) referred to as “hard” GMM. Assuming that the dataset $(\mathbf{x}_{1},\cdots,\mathbf{x}_{N})$ is generated by $K$ independent multivariate normal distributions, we denote the mean vector and covariance matrix of the $k$-th normal distribution as $\bm{\mu}_{k}$ and $\mathbf{\Sigma}_{k}$, respectively. The standard GMM assumes that data points are generated from multiple normal distributions based on a certain probability distribution, with the following likelihood
@@ III-B Statistical Interpretation of Smooth K-Means
where $p_{kn}$ is the probability that the $n$-th data point is generated from the $k$-th normal distribution. The hard GMM, on the other hand, assumes that data points can only be generated from a single normal distribution. Accordingly, for all $n$, $p_{kn}=1$ for one $k\in\{1,\cdots,N\}$ and $p_{in}=0$ if $i\neq k$. We further assume that all normal distributions have the same covariance matrix, specifically an identity matrix. Under these assumptions, the MLEs of $\{p_{kn}\}$ are $\hat{p}_{kn}=1$ if $k=\arg\min_{i\in\{1,\cdots,K\}}\|\mathbf{x}_{n}-\bm{\mu}_{i}\|_{2}^{2}$, and $\hat{p}_{kn}=0$ otherwise. By substituting $\mathbf{\Sigma}_{k}=\mathbf{I}$ and $\hat{p}_{kn}$ into (14) and taking the logarithmic value, we obtain the log-likelihood:
@@
While the hard GMM model is more computationally efficient than the standard GMM, it only considers the impact of a data point on its closest centroid. However, as seen in the success of FKM, it can be advantageous to consider the effect of data points on all centroids for certain real datasets. This can be accomplished by SKM in a parameter-free manner, which applies a smooth function to WCSS. Different smoothing methods result in unique impact patterns, allowing for the capture of various data structures. In the following sections, we will explore three popular smooth minimum functions and their impact on clustering.

### v2

@@ Abstract
Traditional centroid-based clustering algorithms, such as hard K-means (HKM, or Lloyd’s algorithm) and fuzzy K-means (FKM, or Bezdek’s algorithm), display degraded performance when true underlying groups of data have varying sizes (i.e., imbalanced data). This paper introduces equilibrium K-means (EKM), a novel fuzzy clustering algorithm that has the robustness to imbalanced data by preventing centroids from crowding together in the center of large clusters. EKM is simple, alternating between two steps; fast, with the same time and space complexity as FKM; and scalable to large datasets. We evaluate the performance of EKM on two synthetic and ten real datasets, comparing it to other centroid-based algorithms, including HKM, FKM, maximum-entropy fuzzy clustering (MEFC), two FKM variations designed for imbalanced data, and the Gaussian mixture model. The results show that EKM performs competitively on balanced data and significantly outperforms other algorithms on imbalanced data. Deep clustering experiments on the MNIST dataset demonstrate the significance of making representation have an EKM-friendly structure when dealing with imbalanced data; In comparison to deep clustering with HKM, deep clustering with EKM obtains a more discriminative representation and a 35% improvement in clustering accuracy. Additionally, we reformulate HKM, FKM, MEFC, and EKM in a general form of gradient descent, where fuzziness is introduced differently and more simply than in Bezdek’s work, and demonstrate how the general form facilitates a uniform study of KM algorithms.
@@
K-means, fuzzy clustering, imbalance learning, deep clustering
@@ I Introduction
Imbalanced data refers to the true underlying groups of data having different sizes, which is common in datasets of medical diagnosis, fraud detection, and anomaly detection. Imbalanced data poses a challenge for learning algorithms because these algorithms tend to be biased towards the majority group [1]. While there is a considerable amount of research on supervised learning (e.g., classification) from imbalanced data [2, 3, 4], unsupervised learning has not been as thoroughly explored, because the unknown cluster sizes make the task more difficult [5]. Methods like resampling and boosting frequently used in supervised learning cannot be applied in unsupervised learning.
Clustering is an important unsupervised learning task involving grouping data into clusters based on similarity. K-means (KM) is the most popular clustering technique, valued for its simplicity, scalability, and effectiveness with real datasets. It can also be used as an initialization method for more advanced clustering techniques, such as the Gaussian mixture model (GMM) [6, 7]. KM starts with an initial set of centroids (cluster centers) and iteratively refines them to increase cluster compactness. The hard KM (HKM, or Lloyd’s algorithm) [8, 9] and fuzzy KM (FKM, or Bezdek’s algorithm) [10] are the two most representative KM algorithms. HKM assigns a data point to only one cluster, while FKM assigns a data point to multiple clusters with varying degrees of membership.
Research on HKM and FKM has been ongoing. For example, the possibilistic FKM (PFKM) [11] was proposed to address the sensitivity of FKM to noise and outliers. FKM-$\sigma$ [12] was proposed to improve FKM’s performance on data points with uneven variations or non-spherical shapes in individual clusters. Fuzzy local information K-means (FLIKM) [13] was designed to promote FKM’s performance in image segmentation. Deep clustering, a technique that combines deep neural networks (DNNs) with HKM and FKM, was proposed to cluster high-dimensional data [14, 15, 16, 17].
Although many variations based on HKM and FKM have been developed to deal with different situations, most of them display degraded effectiveness in cases of imbalanced data. This is due to the so-called “uniform effect” that causes the clusters generated by these algorithms to have similar sizes even when input data has highly varying group sizes [18]. An illustration of the uniform effect of HKM and FKM is given in Fig. 1 where we can observe that the centroids of HKM and FKM crowd together in the large true cluster.
Figure 1: Clustering results of a highly imbalanced dataset. (a) Ground truth. The colors represent the labels of the data points. (b) Clustering by hard K-means. (c) Clustering by fuzzy K-means. (d) Clustering by the proposed equilibrium K-means. They are all two-step alternating algorithms that iteratively compute centroids.
@@ I-A Existing Efforts to Overcome Uniform Effect
To the best of our knowledge, there are two popular methods to overcome the uniform effect in the field of fuzzy systems. The first method is to introduce more weight on the data points in small clusters at each updating iteration, biasing learning towards them, like a modified FKM called cluster-size insensitive FKM (csiFKM) developed by Noordam et al. [19]. Later, Lin et al. [20] proposed a size-insensitive integrity-based FKM (siibFKM) based on csiFKM to reduce the sensitivity of csiFKM to the distance between adjacent clusters. However, weighting based on cluster size inadvertently increases the influence of outliers, making these algorithms sensitive to noise [21].
The second method is called multiprototype clustering. This method first groups data into multiple subclusters with similar sizes and the final clusters are obtained by merging adjacent subclusters. Liang et al. [22] proposed a multiprototype clustering algorithm that employs FKM to generate subclusters. Later, Lu et al. [5] proposed a self-adaptive multiprototype clustering algorithm that automatically adjusts the number of subclusters. However, multiprototype clustering algorithms have a complex process and high time complexity of $O(N^{2})$, where $N$ is the number of data points in the dataset. Thus, they are computationally expensive for large datasets. We should additionally mention that Zeng et al. [23] recently proposed a soft multiprototype clustering algorithm with time complexity linear to $N$. However, their clustering process remains complex and is aimed at clustering high-dimensional and complex-structured data rather than imbalanced data.
@@ I-B Our Contributions
In this paper, we propose equilibrium K-means (EKM) to address the issue of imbalanced data clustering and the aforementioned algorithms’ limitations. Our contributions can be summarized in three main aspects.
First, we reformulate HKM, FKM, the maximum-entropy fuzzy clustering (MEFC) [24, 25], and EKM in a general form of gradient descent. We show that these algorithms aim to optimize different approximations of the same objective. This general form facilitates the uniform study of KM algorithms.
Second, we develop EKM based on the first contribution. EKM belongs to the family of fuzzy clustering and membership defined in EKM has a clear physical meaning. Repulsive forces appear among centroids of EKM, successfully reducing the uniform effect by preventing centroids from crowding together in a large cluster (see Fig. 1d for an example). Therefore, EKM is effective in dealing with imbalanced data and is not sensitive to noise. Similar to HKM and FKM, EKM is a two-step alternating algorithm that iteratively computes centroids. In addition to having the same time ($O(N)$) and space complexity as FKM, EKM has a batch-learning version that can be applied to large datasets. We demonstrate the effectiveness of EKM on balanced and imbalanced data by conducting numerical experiments on two synthetic and ten real datasets from different domains. Due to space limitations, quantitative performance metrics are provided in the supplementary material. The scatter plots of partial clustering results we provide in the paper are sufficient to demonstrate the superiority of EKM on imbalanced data.
Finally, we investigate the joint learning of DNNs and EKM. We find that mapping high-dimensional data via DNNs to an EKM-friendly space can result in more discriminative low-dimensional representation than mapping to an HKM-friendly space. When tested on an imbalanced dataset derived from MNIST, joint learning of DNNs and EKM improves clustering accuracy by 35% compared to joint learning of DNNs and HKM.
@@ I-C Organization
We introduce HKM, FKM, and MEFC in Section II. In Section III, we derive their general form and the proposed EKM. The properties of EKM are studied in Section IV. We evaluate the performance of EKM on classic clustering tasks in Section V and on deep clustering in Section VI. Finally, we conclude in Section VII.
@@ II-A The Hard K-Means Algorithm (Lloyd’s Algorithm)
KM aims to partition $N$ data points into $K$ clusters, minimizing the sum of the variances within each cluster. Mathematically, the objective can be expressed as:
@@
where $\mathbb{S}_{k}$ represents the set of data points in the $k$-th cluster, $\mathbf{x}\in\mathbb{S}_{k}$ denotes a data point belonging to $\mathbb{S}_{k}$, $\bm{\mu}_{k}$ is the centroid of $\mathbb{S}_{k}$, expressed by
@@
$|\mathbb{S}_{k}|$ signifies the size of $\mathbb{S}_{k}$, and $\|\cdot\|_{2}$ is the $l_{2}$ norm. This optimization problem is NP-hard [26] and is commonly solved heuristically by Lloyd‘s algorithm [9]. Given $K$ points ($\mathbf{c}_{1}^{(1)},\cdots,\mathbf{c}_{K}^{(1)})$ as initial centroids, Lloyd’s algorithm alternates between two steps:
@@ II-B The Uniform Effect of K-Means
The uniform effect refers to the propensity to generate clusters of similar sizes, which is implicitly implied in the objective (1) of KM. For simplicity, kindly consider a case with two clusters (i.e., $K=2$). Minimizing the objective function (1) is equivalent to maximizing the following objective function [27]:
@@ II-C The Fuzzy K-Means Algorithm (Bezdek’s Algorithm)
$\min_{\mathbf{c}_{1},\cdots,\mathbf{c}_{K},\{u_{kn}\}}$ $\sum_{n=1}^{N}\sum_{k=1}^{K}(u_{kn})^{m}\|\mathbf{x}_{n}-\mathbf{c}_{k}\|_{2}^{2}$ (6)
@@
    Calculate the membership value of the $n$-th data point belonging to the $k$-th cluster:
@@
    Recalculate the weighted centroid of the $k$-th cluster by:
@@
The time complexity of one iteration of the above two steps is $O(NK^{2})$. The higher time complexity of FKM than HKM is due to the extra membership calculation.
@@ II-D Maximum-Entropy Fuzzy Clustering
$\min_{\mathbf{c}_{1},\cdots,\mathbf{c}_{K},\{u_{kn}\}}$ $\eta\sum_{n=1}^{N}\sum_{k=1}^{K}u_{kn}\ln{u_{kn}}$ (9)
@@
$\sum_{k=1}^{K}u_{kn}=1\,\forall n,$
where $\eta\in(0,1)$ is a hyperparameter, controlling the transition from maximization of the entropy to the minimization of centroid-data distances. MEFC is similar to FKM but with a different definition of membership:
$u^{(\tau)}_{kn}=\frac{\exp{(-\lambda\|\mathbf{x}_{n}-\mathbf{c}_{k}^{(\tau)}\|_{2}^{2})}}{\sum_{i=1}^{K}\exp{(-\lambda\|\mathbf{x}_{n}-\mathbf{c}_{i}^{(\tau)}\|_{2}^{2})}},$ (10)
@@ III-A Objective of Smooth K-Means
We propose a novel framework called smooth K-means (SKM) and demonstrate that the three KM algorithms introduced in the previous section are special cases of SKM.
Denote the squared Euclidean distance [^1] Other distance metrics, such as the absolute difference and the angle between points, can also be used to define $d_{kn}$. Considering the derivation process is the same, we use the Euclidean distance in the paper for simplicity. between the $k$-th centroid and the $n$-th data point as
@@
and define the within-cluster sum of squares (WCSS) as the sum of squared Euclidean distances between data points and their nearest centroids, i.e.,
@@
where $h(d_{1n},\cdots,d_{Kn})$ is a smooth approximation to $\min(d_{1n},\cdots,d_{Kn})$ and is referred to as the smooth minimum function. Below we reveal the relationship between the SKM and the GMM.
III-B The Relationship Between SKM and GMM
The centroids obtained by minimizing WCSS are maximum likelihood estimators (MLEs) of parameters of “hard” GMM. The derivation is as follows. Assuming that the dataset $(\mathbf{x}_{1},\cdots,\mathbf{x}_{N})$ is sampled from $K$ independent multivariate normal distributions. We denote the mean vector and covariance matrix of the $k$-th normal distribution as $\bm{\mu}_{k}$ and $\mathbf{\Sigma}_{k}$, respectively. The standard GMM has the following well-known likelihood
@@ III-B The Relationship Between SKM and GMM
where $p_{kn}$ is the probability that the $n$-th data point is generated from the $k$-th normal distribution. The standard GMM assumes that data points are generated from multiple normal distributions based on a certain probability distribution. On the other hand, the hard GMM assumes that a data point is generated from a single normal distribution. Accordingly, for all $n$, $p_{kn}=1$ for one $k\in\{1,\cdots,N\}$ and $p_{in}=0$ if $i\neq k$. We further assume that all normal distributions have the same covariance matrix, specifically an identity matrix. Under these assumptions, the MLEs of $\{p_{kn}\}$ are $\hat{p}_{kn}=1$ if $k=\arg\min_{i\in\{1,\cdots,K\}}\|\mathbf{x}_{n}-\bm{\mu}_{i}\|_{2}^{2}$, and $\hat{p}_{kn}=0$ otherwise. By substituting $\mathbf{\Sigma}_{k}=\mathbf{I}$ and $\hat{p}_{kn}$ into (14) and taking the logarithmic value, we obtain the log-likelihood:
@@
The hard GMM model simplifies the standard GMM by only considering the impact of a data point on its closest centroid. However, as seen in the success of FKM, it can be more advantageous to consider the impact of a data point on all centroids. This can be accomplished in SKM by applying an approximation function to smooth WCSS. Hence, SKM can be viewed as a model between hard and standard GMM. In the following sections, we will introduce three common smooth minimum functions and explore the resulting distinct clustering algorithms.